4 Decmeber, 2016

Had visitors at the observatory in the evening. A young and a very enthusiastic child, who is just 5, saw the planets and the moon with the 127 refractor telescope. 

Main observatory was closed. 

3 December, 2016

17:30

In the early evening, two students doing A levels, came to the observatory and i showed them the observatory. Thy are very interested in learning how the telescopes at the observatory works and what can we do with this equipment.

We used Sky Watcher 10 inch Dobsonian telescope to first see the cresent moon. They became very excited! Then we saw nearby and very brilliant Planet Venus. Mars was next and the last was Albireo, the colorful double star.

We also grabbed some quick pictures from the phone camera. 

2 December, 2016

Lahore Astronomical Society had its monthly meeting at the observatory. Telescopes were closed.

30 November, 2016

19:30

Sky is partly cloudy.. and hazy too.

20:45

I have installed the newly arrived GPS receiver in the mount and it is working all fine!

21:37

The humidity is very high.. 79% right now in the observatory.. not good.

22:11

Sky is getting worst here.. I cannot attempt any deep sky objects tonight. I am going to attempt to collect some astrometric data.

22:16

TCF is not focusing.. the focus point has moved beyond 7000.. beyond the limit of TCF. I now need to change the focus manually and readjust few more things.

22:58

Alright.. i have loosened the mirror holding of C14 telescope, changed the TCF value from 7000 to 2500, have manually refocused the scope with the help of Bahtinov mask, collimated the optics again, refocused the system with the automated TCF Focuser and now the star is perfectly focused!

23:07

The humidity level in the observatory room has gone upto 94%!! Things are dripping, literally!

I have turned off everything and closed the roof.

29 November, 2016

18:45

Roof is opened. Sky is not that good as it was yesterday and the day before it. I could barely see 3.5 magnitude star, which was Achird in Cassiopeiae constellation.

Waiting for the system to cool down.

The 'Cloud Sensor' was reading the clarity aroun 25 these past two nights; tonight it is 17.. bad news indeed!

19:16

I live very near to a big market where they throw so much light on the sky that the eastern part is heavily light polluted and this it the hemisphere where i almost always image. Further there is a huge screen they have placed, which is just half a mile from my observatory. Cannot do anything about it.

20:03

Now taking luminance frames of NGC 891 galaxy. Since the sky is not so good tonight, it is looking very dim on my screen.

21:23

Experimented on focusing.. going to write a short blog on it.

Photons from NGC 891 still coming through.

21:54

This i just wrote: http://umairasim.com/blog/2016/11/29/temperature-effect-on-focusing

23:11

Took a quick photometric V filter data of V0725 Tau, analyises and submitted to AAVSO.

23:41

Observatory is closed now. 

 

28 November, 2016

18:30

Opened the roof.. tonight the sky is also very promising.

18:53

Need to take some flats for luminance filter.

19:12

I found a wooden circular thing they use for art in a stationary store. Immediately realized that i can use it for taking flats. Just going to experiment on this.

19:23

Success!! It worked perfectly :) Now on to flat images.

21:45

Adding image rotator has made my life so easy and so difficult at the same time. With camera angle rotation, light cone is also rotated. So i need lots and lots of flat frames. Doing that right now.

10:07

Stopped the flats thing and on the one of the asteroids i am observing these days.

10:27

Imaging has begun.. I will at least take an hour worth of data of this asteroid for a good astrometry.

12:02

Collected 43 images, each of 90 seconds.. I will do the astrometric analysis later.

03:47

Got images of a beautiful part of Orion Nebula.. an hour worth of data. Imaged another asteroid.

Having difficulties with the TCF focuser.. i have reached the end of its position. Probably tomorrow i will change the focuser of the OTA and then there are many steps before i can restarty imaging routine again.

 04:24

Observatory is closed now. 

 

27 November, 2016

19:15

Slided the roof out. Sky is very good at the moment.

20:05

The Sky X is showing NGC 891 to be at a very good position and this is an edge on galaxy and one my my most favourite one.

Lets take some luminance data of this galaxy, before i start doing some serious work :)

20:29

Focused the system and now a magnitude 10.37 star is guiding the mount at 2.2 Hz, which is very good!

First image came nice.. this is such a beautiful galaxy.

21:03

27 images have been downloaded, each of 60 seconds. I have decided to keep single exposures short, even though guiding is at 2.2 Hz. This will keep me safe, if anything bad happens with the guiding.

The USB camera cable just got disconnected! After reconnecting, i saw the image was right in the center. More photons from NGC 891 are coming through and being captured at my camera chip.

21:17

Just realized, i have no started periodic error connection on my mount, it is working now.

22:29

Still getting more of NGC 891. This galaxy has passed the maridian line but my mount can track some a few degrees before it needs to flip and change side.

01:30

Got some luminance data on Eskimo Nebula and an interacting galaxies' pair. 

Also some astrometry of an asteroid. 

Observatory is closed now. 

26 November, 2016

18:30

By the time night started, it is all crisp and clear. Virtually the haze is gone. 

19:30

Patches of clouds came from the west. These days i first go to Capella, after i start my system, to check the pointing accuracy. I could see Capella clearly.. now it is behind the clouds.

Camera is cooling down and it is at -18C right now. I image at -30 C.

20:36

The telescope is at the target right now but clouds have covered the two third of the sky right now. Only the west side is clear. I hope these clouds will pass in about 15 minutes.

I and the telescope are all ready for this Asteroid tracking.

21:02

Mount tracking is not being good to me. Somehow the mount keeps pushing the scope towards east. I have yet to figure out why this is happening.. may be some backlash issue.

21:10

Problem solved! Auto guiding is on and the guide star is dead center now.. happy me :)

23:40

I have tracked three asteroids so far. Analyzing the data now.

V0725 Tau is the next target. V band is done, B has just started.

01:23

have finished tri band BVI data aqcuisition, processed these images and aqcuired the magnitude of V0725 Tau star. Submitted to AAVSO.

Also got some images of LL Ori in luminance filter. Seeing became a bit bad.. the guide star was jumping all around even at 16.7 Hz of autogiiding. 

02:40

Observatory is closed now.. going to take a good night sleep! 

 

24 November, 2016

19:30

Sky was cloudy all day but became better in the evening. I have just come from an outreach session for young children, which just gave me another shot of 'redbull' to continue and expand my outreach programs.

20:20

The sky is all clear.. at least for now. I have opened the observatory's roof and the sequence of operation has started.

20:53

System is all ready, focused, on target, guide star found and being guided right now. Imaging has begun!

23:27

A good tracking of an asteroid has been performed for about an hour and a half.

Now on to photometry stuff!

00:10

I Filter covered.. now switching to V. Autoguiding is been amazing!

Sky is still clear, no clouds.. though a bit hazy.

00:41

I and V reports submitted.. data looks quite good.

B is coming..

01:15

Submitted all thre band to AAVSO. Closing the observatory now.

23 November, 2016

19:28

The sky is bad.. Cloud sensor is indicating 'very cloudy' sign on the screen. Observatory is closed.

It remained the same the whole night and i could not open the observatory.

22 November, 2016

18:06

Sky is partly cloudy and the haze is all over. Just opened the roof of the observatory.

18:15

Adjusted some wired. Checked the water reservoir.

18:29

Started taking dark frames of 600 seconds.

Waiting for my target asteroids to move up in the sky where they can be imaged.

20:47

Me and my memory.. I thought i am taking darks and just noticed, i had yet to push the button! Precious time has been wasted.

21:12

Sky is so terribly bad! I could barely see one of the brightest star Capella. Still trying to get any image of an asteroid.

22:51

78 images are in.. more are coming.. Asteroid 468 Lina is looking alright.

23:34

Tried other asteroids but the stars in their field of view are too dim and no astrometry is possible for these rocks. Sticking by the same asteroid.

114 minutes of continuous tracking of this asteroid.

12:28

468 Lina has crossed the meridian line. Next target Asteroid 648 Pippa is at 54 degrees altitude and in a good amount of surrounding stars in the FOV with a nice star in my guiding chip. Pyxis Rotator angle is 22 degrees.

Refocused the telescope, slewed back to the asteroid, autoguiding started at 0.6 Hz, exposure for the main chip is set to 60 seconds, made the appropriate folders and imaging has started.

03:03

Posted an animation of 468 Lina on the blog. With this sky conditions, no photometry will be wise. Astrometry is the only way to go in my opinion.

Observatory is closed now. Moon is all blurred in the sky right now.

21 November, 2016

17:00

Came to the observatory at maghrib (sunset), getting everything ready for tonight's imaging. Have an astrometry of some asteroids in mind. Let's see if i could find some decent targets.

18:11

arrgghhh.. clouds came! Closed the roof immediately.

18:46

The sky is completely covered with clouds now.

I have started taking Dark frames of 600 seconds each.

19:32

Clouds are getting thicker.. nothing up there i can do now. Time to play with the images i already have.. and i have plenty to survive me weeks.

20:48

Attempting to make an animation from my last lunar images.

Animation is successful.

23:10

Sky is clear now but the haze is there. I am powering up the observatory now.

00:02

Found a good target which is bright enough to be imaged in this hazy sky.. an asteroid. It took me about 40 minutes to move the telescope to the correct position. Plate solving was being very difficult because stars' light is not reaching to the CCD.

01:07

Got a good amount of data of an asteroid. 

Now a patch of cloud is right at the zenith.. waiting for it to move away so i can resume another asteroid imaging.

01:30

The sky is clear again.. Now imaging 468 Pippa in luminance filter.

02:19

Cloud cover.. observatory is closed now.

19 November, 2016

21:20

Opened the roof and the SOP of the observatory began, in which number of steps are involved before i can actually start imaging runs.

22:52

First image of V0725 Tau is received from the camera in I Band.

11:40

100 images have been collected in I band.

01:51

100 images of V and B band have been acquired now. The last image received at 01:51.

02:15

 Rarely i have attempted moon with C14 telescope and SBIG ST9XE camera. I have a very sensitive CCD and moon is always way more brighter than my chip can record. But this night, i am going to work seriously on my setup to see if i can actually get a nice lunar image.

I will write a blog with all the details of the imaging process i used. In short, i used an infrared filter and got many frames of 'Sea of Serenity' on the moon. The images were from 02:15 to 02:30.

03:10

Next run was from 03:10 to 04:24 and during this time i acquired 2, 569 images of the same area. 

05:10

Submitted the data to AAVSO and closed the roof.. time to head home!

18 November, 2016

21:12

Observatory is being ready for tonight's observations.

Dark frames library is being updated in the past few days. Next group is of 240 seconds Dark frames.

Some guests came to visit the observatory.. They were shown the telescope fully engaged in doing, what it is being told to :)

22:47

22 Dark frames of 240 seconds have been collected.

23:11

Slewing to V0725 Tau star in Taurus constellation. I band will be the first to capture.

00:29

160 frames have been acquired. Maxim DL, as always, is being used to 'clean' all images.

00:54

V0725 Tau is reported as 7.773 magnitude in I band to AAVSO.

Now on to V band.

01:50

V data reduced and analysed. The reading of 8.741 magnitude is submitted to AAVSO.

Onto B band now!

17 November, 2016

21:56

Dark frames of 30 Seconds are being exposed.

22:37

Slewed the telescope to V0725 Tau star. Since the star is comparatively low in the sky, it makes sense to start taking the I band data.

Focused the system to a near by 4.5 magnitude star.

00:16

Got 130 frames of I band with 15 seconds of exposure. Tonight i want to add V band as well.

Checked the exposure time of V0725 Tau star and it turns out that it is dim (as stars usually are) in V band. A 30 seconds of exposure gives me a peak pixel value of around 50k. Generally this is not a good value for photometry data but fortunately i have a NABG CCD so for me, it is all fine!

00:36

Guide star acquired and now my mount is being guided at 1 Hz, which is very nice indeed. The data is being collected now in V band.

01:14

After V data, V0725 Tau star is still in the eastern hemisphere. I am going to attempt some B band too.

As expected The star is dimmest in B band. A 60 seconds of exposure will do. Autoguiding is set to 0.4 Hz. 

01:22

Just Hit the "Take Series" knob in the Sky X Pro.

02:44

BVI done. Now time to take some flats. Arranged the scope for my flat acquisition setup. I would take at least 100 Flats for B and V filters. I already have a library for I filter.

04:58

Data reduced and all three BVI photometry bands are analysed. Submitted to AAVSO.

 

 

 

16 November, 2016

18:50

The call for prayer of Isha marks the end of the twilight and the start of night. Just opened the roof of the observatory.

Tonight's targets are: V0725 Tau Photometric Data and some dark frames at -30C. CCD is at -25C right now.. still cooling down.

Water reservoir is at 19C ; i have set the cooling point to 13C.

21:10

50 Dark Frames of 120 seconds each have been acquired. 

V0725 Tau is still at 27 degrees altitude (very low for any scientific data acquisition) in the eastern horizon.

00:12

Focused the telescope with photometric I filter. Autoguiding at a 10th magnitude star at 0.5 Hz with the combination of Adaptive Optics and Mount autoguiding capability.

90 Seconds of exposure still saturated the star.

00:20

Changed it to 30 seconds now.. nope, still too much light! Lowered it to 15 seconds now and voila! a very nice 40K counts.

Acquisition begins now!

03:15

Have got 162 images in total. Now taking 15 sec Darks for some time now.

04:51

Reduced the data and performed photometry analysis in Maxim DL. As it turned out V0725 Tau star has right now 7.855 magnitude in I band. Submitted the data to AAVSO.

15 November, 2016

21:30

Opened the roof of the observatory around 10pm; telescope and the mount will cool down to the ambient temperature now.

An AAVSO alert just came in: A request for multiband photometry of a star in Taurus: "V0725 Tau".

All the equipment got connected easily, except the Pyxis 3 Rotator.. tried everything, unplugging and plugging back, reinstalling the drivers, restarting the Sky X Pro software and re starting windows. Nothing worked! Finally found out, the port has mysteriously been changed. It is working fine now. It just wasted my one and half an hour.

Camera cooling point is changed tonight. So far i had adjusted it at -20C.. from tonight, it is at -30C!!

An hour was left before the star will cross meridian line.

Found the star chart from AAVSO.. I do have at least one reference star in my field of view.

AAVSO has requested for photometric V filter data as a priority but since moon is so close to the star, i have decided to use photometric I filter, which will cut down a lot of sky background light.

Slewed the telescope to a nearby 5 magnitude star, focused the telescope and returned to my target area.

A guide star of of magnitude ~10.5 is available. After calibrating the autoguider setup with the Pyxis 3 Rotator's Position Angle of 144, the guide star is now being tracked at 0.5 Hz.

Made the folders for the images, adjusted the camera settings with the exposure of 180 seconds and hit the 'Take Series' button in Sky X Pro.

A few images came bad.. it seems the fog has started to build up in the CCD chamber because of very low temperature settings. I would have to bake the camera desiccant very soon.

On 6th image the CCD chip is now doing fine... Imaging, continues..

02:30

Ah.. unfortunately.. me being careless, did not test  the exposure time of the single images. 180 seconds turned out to be too bright for the target star and i somewhat saturated it. In simple words, all the data so far taken is useless now.

I have changed the temperature of the CCD now.. so i need to take dark frames. Doing that now. 30 Seconds darks are done.. 60 seconds in progress.

05:00

30, 60 and 90 seconds dark frames have been recorded. Roof closed now!